OpenStudy (anonymous):

Verify/establish the identity: [(2sin^2 theta-1)^2/(sin^4 theta-cos^4 theta)=1-cos^2 theta I feel like I should work with the LHS in verifying the RHS.

zepdrix (zepdrix):

\[\Large\rm \frac{(2\sin^2\theta-1)^2}{(\sin^4\theta-\cos^4\theta)}\]Ooo this is an interesting problem :)

zepdrix (zepdrix):

Bottom we have the `difference of squares`, it might not be super easy to notice at first though.

zepdrix (zepdrix):

Remember how to factor difference of squares into conjugates? Looks like this: \[\Large\rm a^4-b^4=(a^2)^2-(b^2)^2=(a^2-b^2)(a^2+b^2)\]

zepdrix (zepdrix):

After that, I would probably apply the `Cosine Double Angle Formula` to top and bottom. Need to see a few more steps? :o

zepdrix (zepdrix):

I might be wrong about the Double Angle, I'm trying to do too many steps in my head D:

OpenStudy (anonymous):

If that's okay. I just learned about Double angle and half-angle formulas today and I'm still fairly uncertain. Thank you for helping me again!

zepdrix (zepdrix):

Ok I applied 2 steps here, \[\Large\rm =\frac{(1-2\sin^2\theta)^2}{(\sin^2\theta-\cos^2\theta)(\color{royalblue}{\sin^2\theta+\cos^2\theta})}\]In the numerator I factored a -1 out of each term (it get's squared when coming out of the brackets and becomes positive). In the denominator we applied the difference of squares formula. Notice anything interesting about the blue portion?

zepdrix (zepdrix):

It seems like this is going to end up simplifying to \(\Large\rm 1-2\cos^2\theta\) Are you sure you wrote it down correctly? :o Again I'm doing these steps in my head >.< so I could be wrong...

OpenStudy (anonymous):

Yes, the change in sign. Is that due to the difference of squares or is that the cosine double angle formula?

zepdrix (zepdrix):

Hmm? Are you talking about the numerator or denominator step?

zepdrix (zepdrix):

Oh I asked about the blue part :) lol. Yes that's due to the difference of squares formula.

OpenStudy (anonymous):

Oh sorry, I forgot to type the two. I am so terribly sorry!

OpenStudy (anonymous):

The denominator.

zepdrix (zepdrix):

lolol you apologize too much ^^ simmer down ms Alyssa!! so proper! lol

OpenStudy (anonymous):

Okay, just making sure.

OpenStudy (anonymous):

Then how do you implement the cosine double angle formula from here, if I may ask.

zepdrix (zepdrix):

That's not what I wanted you to notice about it though!! :) The blue part is our Pythagorean Identity for sine and cosine yes? That simplifies down REALLY nice.

OpenStudy (anonymous):

Ohhhh I hadn't noticed that! I'm sorry, I'm really focused on the angle formulas.

zepdrix (zepdrix):

I guess you don't have to apply the Cosine Double Angle Formula if you're uncomfortable with it. It's just that we want to try and make these orange parts the same,\[\Large\rm =\frac{(\color{orangered}{1-2\sin^2\theta})^2}{(\color{orangered}{\sin^2\theta-\cos^2\theta})(\color{royalblue}{\sin^2\theta+\cos^2\theta})}\]

zepdrix (zepdrix):

We can do that by simply using the Pythagorean Identity and changing our cos^2(theta) to 1-sin^2theta

OpenStudy (anonymous):

Thank you! I'm sorry that I didn't see it before!

zepdrix (zepdrix):

LOL there you go again XD

zepdrix (zepdrix):

\[\Large\rm =\frac{(\color{orangered}{1-2\sin^2\theta})^2}{(\color{orangered}{\sin^2\theta-(1-\sin^2\theta)})(\color{royalblue}{1})}\]

zepdrix (zepdrix):

Which leads us to,\[\Large\rm =\frac{(\color{orangered}{1-2\sin^2\theta})^2}{(\color{orangered}{2\sin^2\theta-1})}\]AHHH I shoulda left the top alone.... Remember when I factored out that negative earlier? blah!

zepdrix (zepdrix):

So put the negative back in XD\[\Large\rm =\frac{(\color{orangered}{2\sin^2\theta-1})^2}{(\color{orangered}{2\sin^2\theta-1})}\]

OpenStudy (anonymous):

Yes, I remember. Don't worry, I'm just grateful that you're guiding me through this problem.

zepdrix (zepdrix):

So if we haven't made any mistakes up to this point, it looks like we get a nice cancellation,\[\Large\rm =\frac{(2\sin^2\theta-1)^{\cancel2}}{\cancel{(2\sin^2\theta-1)}}\]

zepdrix (zepdrix):

But as I mentioned before, this is not going to simplify to \(\Large\rm 1-\cos^2\theta\). So I hope the problem was written down incorrectly D:

OpenStudy (anonymous):

Yes, I'm sorry again. It's supposed to simplify to 1-2cos^2 theta

OpenStudy (anonymous):

Thank you so much for being so patient with me.

zepdrix (zepdrix):

Ah ok great :) that looks better.

zepdrix (zepdrix):

So you're left with,\[\Large\rm 2\color{#CC0033}{\sin^2\theta}-1\]And you would just apply your Pythagorean Identity on this sin^2(theta) a final time to get your answer. Any of those steps too confusing? :o

zepdrix (zepdrix):

Like this, \[\Large\rm =2(\color{#CC0033}{1-\cos^2\theta})-1\](Just in case there was any confusion with that step >.< )

OpenStudy (anonymous):

Not anymore. Once I saw where you redistributed your factor. I really cannot thank you enough for all your help.

zepdrix (zepdrix):

This problem is a bit of a doozy! Lot of little things to remember, most importantly: Difference of Squares.

zepdrix (zepdrix):

Aw great c: Glad to help!

OpenStudy (anonymous):

Yeah, sometimes it's difficult to recognize where I need to apply difference of squares, but it comes with practice, right. :) Thank you again! For being such a kind and thoughtful instructor!

zepdrix (zepdrix):

G'day M'lady ^^/ See you at the next problem perhaps. hehe

OpenStudy (anonymous):

I hope you have a wonderful day as well! I would be overjoyed to have you guide me again!